A ceiling fan is turned on and a net torque of 1.6 N · m is applied to the blades. The blades have a total moment of inertia of 0.24 kg · m2. What is the angular acceleration of the blades?
Answer #1

Torque \tau = 1.6 N

moment of inertia I = 0.24 Kg-m2

From the formula   \tau = I .\alpha

hence angular acceleration \alpha =  \tau / I

= 1.6 / 0.24

= 6.67 rad/s2